passes through (-3,-1), parallel to the line that passes through (3,3) and (0,6)
Answer by swincher4391(1107)

Line that passes through (3,3) (0,6).
Step one: Find slope
(6-3)/(0-3) = 3/-3 = -1
Create a line that is parallel to one with slope -1 and goes through point (-3,-1). This screams point slope form. Where our slope is -1 (cause parallel lines have the same slope) and our point is (-3,-1)
(y+1) = -1(x+3)
y +1 = -x -3
y = -x -4